OpenStudy (anonymous):

Use the distributive property to simplify the expression below: (12x-10)/(-2) −6x + 5 12x + 5 −6x − 5 6x + 5

OpenStudy (anonymous):

the second one

OpenStudy (ajprincess):

\(\Large\frac{12x-10}{-2}\) \(=\Large\frac{12x}{-2}-\frac{10}{-2}\) \(=-6x-(-5)\) \(=-6x+5\) Is it clear?